# https://www.projectaria.com/datasets/nymeria

### v1

@@ 6 Discussions and Conclusions
The appendix provides further details about the dataset and algorithms. The structure is as follows. Appendix 0.A gives more details about the capture setup. Appendix 0.B provides descriptions about data processing, including a brief summary of Project Aria MPS [7] in Sec. 0.B.1, XSens motion retargetting in Sec. 0.B.2, and motion-language narration interface in Sec. 0.B.3. Appendix 0.C provides additional oversights of dataset construction w.r.t. recording scenarios (Sec. 0.C.1) and locations (Sec. 0.C.2). Appendix 0.D discusses details about the baseline experiments. Appendix 0.E presents the contribution statement and acknowledgement.
@@ 0.B.2 Motion retargeting
Among the 89 body joints, we use one joint to represent the global transformation in the world. 35 joints are designed to represent bones in the human body and they loosely follow the human anatomical structure (e.g. the forearm joint corresponds to the ulna and serves as the twist pivot of the wrist, which is located at the tip of what would be the radius). Fig. 10 (middle, right) overlays the joints on a skeleton model for reference. In addition, we designed 53 "helper" joints for the purpose of improving skinning deformation, as commonly done in computer graphics. These helper joints are driven by the 35 bones with a simple linear relationship. For example, there are five helper joints along each upper and lower limb respectively. Their purpose is to distribute the twist rotation values between the two end joints along the limbs, to avoid the "candy wrapper" artifact of linear blend skinning.
@@ 0.D.1 Baseline methods
Appendix 0.E Contributions and Acknowledgements
0.E.0.1 Contribution.
Lingni Ma led the project, developed the pipeline to construct the dataset, coordinated data collection/processing, and led the baseline evaluations. Yuting Ye developed the solution for human motion retargetting, advised data collection and evaluations. Fangzhou Hong, Vladimir Guozv and Yifeng Jiang, validated the dataset and implemented baselines. Rowan Postyeni supported daily operations, processed XSens motion data and performed quality assessment for XSens motion and narration. Luis Pesqueria served the program manager for data collection and narration. Alexander Gamino was responsible for multi-device tracking. Vijay Baiyya was responsible to Project Aria MPS scaled processing. Hyo Jin Kim supported narration annotations. Kevin Bailey and David Soriano Fosas lead hardware development of miniAria wristband. C. Karen Liu and Ziwei Liu acted as technical advisor to data collection, annotation and baseline evaluations. Jakob Engel, Renzo De Nardi and Richard Newcombe were the senior technical and scientific advisors.
0.E.0.2 Acknowledgement.
We gratefully acknowledge the following colleagues for their valuable discussions and technical support. Genesis Mendoza, Jacob Alibadi, Ivan Soeria-Atmadja, Elena Shchetinina, and Atishi Bali worked on data collection. Yusuf Mansour supported gaze estimation on Project Aria. Ahmed Elabbasy, Guru Somasundaram, Omkar Pakhi, and Nikhil Raina supported EgoBlur as the solution to anonymize video and explored bounding box annotation. Evgeniy Oleinik, Maien Hamed, and Mark Schwesinger supported onboarding Nymeria dataset into Project Aria dataset online browser. Melissa Hebra helped with coordinating narration annotation. Tassos Mourikis, Maurizio Monge, David Caruso, Duncan Frost, and Harry Lanaras provided technical support for SLAM. Daniel DeTone, Dan Barnes, Raul Mur Artal, Thomas Whelan, and Austin Kukay provided valuable discussions on annotating semantic bounding box. Julian Nubert adopted the dataset for early dogfooding. Pedro Cancel Rivera, Gustavo Solaira, Yang Lou, and Yuyang Zou provided support from Project Aria program. Edward Miller served as research program manager. Pierre Moulon provided valuable guidance to opensource codebase. Svetoslav Kolev provided frequent feedbacks. Arjang Talattof supported MPS. Gerard Pons-Moll served as senior advisor. Carl Ren and Mingfei Yan served the role of senior managers.

### v2

@@ Acknowledgements
The appendix provides further details about the dataset and algorithms. The content is structured as follows. Appendix 0.A gives more details about the capture setup. Appendix 0.B provides descriptions about data processing, including a brief summary of Project Aria machine perception service (0.B.1), XSens human model retargetting (0.B.2), and motion-language annotation interface (0.B.3). Appendix 0.C provides additional information of data recording scenarios (0.C.1) and locations (0.C.2). Appendix 0.D presents details about the baseline experiments.
@@ 0.B.2 Motion retargeting
Among the 89 body joints, we use one joint to represent the global transformation in the world. 35 joints are designed to represent bones in the human body and they loosely follow the human anatomical structure (e.g. the forearm joint corresponds to the ulna and serves as the twist pivot of the wrist, which is located at the tip of what would be the radius). Figure 10 (middle, right) overlays the joints on a skeleton model for reference. In addition, we designed 53 "helper" joints for the purpose of improving skinning deformation, as commonly done in computer graphics. These helper joints are driven by the 35 bones with a simple linear relationship. For example, there are five helper joints along each upper and lower limb respectively. Their purpose is to distribute the twist rotation values between the two end joints along the limbs, to avoid the "candy wrapper" artifact of linear blend skinning.